OpenStudy (liv33):

Medal and fan. Which of the following is not true? A. 5-6=-6+5 B. 3(x-5p)=3x+15p C. -8x=x(-8) D. 4(5-2t)=-8t+20

Vocaloid (vocaloid):

any ideas?

OpenStudy (liv33):

A?

Vocaloid (vocaloid):

what does 5 - 6 equal?

OpenStudy (liv33):

-1

Vocaloid (vocaloid):

good, now what is -6 + 5?

OpenStudy (liv33):

-1

OpenStudy (liv33):

ok so it's actually true

Vocaloid (vocaloid):

any other ideas?

Vocaloid (vocaloid):

hint: it involves the distributive property

OpenStudy (liv33):

B

Vocaloid (vocaloid):

good!

OpenStudy (liv33):

Because 3(x-5p)=3x+15p = 3x-15p=3x+15p (different signs) ?

OpenStudy (liv33):

Thanks :)

Vocaloid (vocaloid):

yes, very good